12. A tool rental company determines that it will rent 500 jackhammers per day at a daily rental fee of $30 per jackhammer. For each $1 increase in rental price, 10 fewer jackhammers will be rented. What rental price maximizes revenue?
Answer by Boreal(15235)

the revenue function is (500-10x)(30+x)=15000-300x+500x-10x^2
=-10x^2+200x-15000
this has a maximum at x=-b/2a=-200/-20 or 10
when the price is $40, the number is 400, and the revenue is $16,000
check with $41 and 390 ($15,990) and 39 and 410 (same)